I need help solving the following problem.... 4(x+4)=6(x-2)
the answer must be in simplitied form.

Answer by Alan3354(69443)

Multiply to eliminate the parentheses.
4(x+4) = 4x + 16
6(x-2) = 6x - 12
------------
So 4x + 16 = 6x - 12